OpenStudy (anonymous):

a box measures 1/6ft. long ,1/4ft. wide, 1/3ft. high .Find the volume of the box in inches units .Do not include units in your answer .

OpenStudy (anonymous):

volume = 1/72 cu ft 1 cu ft = 12^3 = 1728 cu ins vol of box = 1728 / 72 = 24